OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <2, -4>, v = <3, -8> *MEDAL WILL BE AWARDED*

geerky42 (geerky42):

Use this formula: \(\vec{u}\cdot \vec{v} = |\vec{u}||\vec{v}|\cos\theta\)

OpenStudy (anonymous):

too late @geerky42

geerky42 (geerky42):

k